# ReCLIP++: Learn to Rectify the Bias of CLIP for Unsupervised Semantic Segmentation

### v1

@@ IV-C Comparison with Conference Version
[…]
@@
In Table III, we verify the effectiveness of each technical improvement of ReCLIP++ beyond ReCLIP [27] (our conference version): 1) we additionally introduce a decoder which takes the rectified logit map and visual feature of CLIP as input and outputs smoother and more contextual masks (denoted as “Decoder”); 2) we design a new strategy which generates more accurate label hypothesis for each image (denoted as “Label Hypothesis”); 3) we optimize the design of bias encoding scheme to independently encode the class-preference and space-preference bias (denoted as “Independent”). In Table III, we observe that when sequentially adding each technical component, the segmentation performance is consistently improved, which verifies the effectiveness of each technical contribution.
@@
Finally, as shown in Table II, with all these technical improvements, ReCLIP++ remakably outperforms ReCLIP. For example, on PASCAL VOC and Cityscapes, ReCLIP++ outperforms ReCLIP by 3.4% and 7.0% respectively. Notably, we observe that only with rectification stage, ReCLIP++ (rec.) achieves comparable or even better performance than ReCLIP (after rectification and distillation stage), which exhibits stronger bias rectification capability of ReCLIP++ compared to ReCLIP. For example, on PASCAL Context, ReCLIP++ (rec.) is comparable to ReCLIP (34.0% vs. 33.8%); on Cityscapes, ReCLIP++ (rec.) outperforms ReCLIP by around 2%.
IV-D Ablation study
TABLE IV: Ablation on whether Query should be learnable. Results after the rectification stage are reported.
Query Reference PASCAL VOC PASCAL Context
[…]
@@ IV-D Ablation study
Should Query be learnable? For text inputs $Q$ and $R$, we only make Reference $R$ learnable while keeping Query $Q$ fixed. In this ablation, we study on whether the prompt of Query should also be learnable. As shown in Table IV, we conduct experiments with learnable and fixed Query respectively. The numbers show that fixing Query obviously outperforms making Query learnable. The reason is when we make Query learnable, it may also implicitly capture bias, confusing the following bias rectification operations (e.g., bias logit subtraction). Therefore, in our framework, we choose to fix $Q$ to make the bias rectification more effective.
@@
Effect of element-wise subtraction on bias rectification. In order to validate the effect of our element-wise subtraction, we conduct experiments in Table V. We compare our subtraction mechanism with an alternative solution: instead of subtracting the bias logit map $M_{b}$ from the Query logit map $M_{q}$, we add $M_{b}$ to $M_{q}$. Comparisons shown in the table verify the effectiveness of our subtraction way. As we aim to remove the bias from the original CLIP, we speculate that the subtraction operation may work as a strong prior which regularizes the training to facilitate meaningful and effective bias modeling.
TABLE V: Effectiveness of element-wise subtraction. Results after the rectification stage are reported for comparison.
Subtraction Addition PASCAL VOC PASCAL Context
[…]
Sensitivity to $t$ and $r$ in image-level multi-label hypothesis generation.
Figure 3: Sensitivity to $t$ and $r$ in image-level multi-label hypothesis generation. (a) Sensitivity to threshold $t$. (b) Sensitivity to $r$. Results after rectification stage are reported.
@@
When generating image-level multi-label hypothesis as discussed in Sec. III-D, $t$ and $r$ need to be determined. We study the sensitivity of our method to these hyper-parameters. As shown in Fig. 3, we observe that the mIoU of our method firstly increases and then decreases as $t$ or $r$ increases, exhibiting a typical bell curve. Such a curve shows the regularization effect of $t$ and $r$ on the training. By replacing the multi-label hypothesis generation solution in ReCLIP++ with that used in ReCLIP (i.e., “w/o. crop” in the figure), we obtain results which are inferior to ReCLIP++ within a vast range of $t$ or $r$. All the results show that the solution adopted in ReCLIP++ is insensitive to the choices of $t$ and $r$ and is superior than that used in ReCLIP.
@@
Effect of Fine-tuning CLIP. We conduct experiments with fine-tuning the image encoder of CLIP in our framework. The results on PASCAL VOC and PASCAL Context are shown in Table VI. We find a consistent performance decrease (to nearly 0 mIoU) compared to freezing CLIP. It is because fine-tuning CLIP on small-scale downstream datasets may destroy the image-text alignment in pre-trained CLIP.
@@
TABLE VI: Effect of Fine-tuning CLIP. The “Fine-tune CLIP” means we fine-tune the image encoder of CLIP.
Setting PASCAL VOC PASCAL Context
Fine-tune CLIP 1.2 0.4
Frozen CLIP 73.5 34.0
@@
Effect of Bias Rectification. In this ablation, we further evaluate the rectification effect on two biases both qualitatively and quantitatively.
@@
(1) Qualitative Results of Bias Rectification. In Fig. 4 (a), we illustrate how we explicitly model and rectify the class-preference bias. As shown in the first column of Fig. 4 (a), the original CLIP (see “Query” mask) tends to misclassify part of a “person” (see the ground-truth mask denoted as “GT”) into a “boat”. Such a mistake is reflected in the comparison between the logit heatmaps for the “person” channel and the “boat” channel: in the area misclassified as “boat”, the logits for “boat” are relatively higher than those for “person”. In contrast, for the bias logit map, the person-channel logits are quite low, while the boat-channel logits are generally very high, especially for the misclassified area. Consequently, via the proposed element-wise logit subtraction operation, we obtain the rectified logit map (see the last column), where the boat channel is largely suppressed. Finally, we obtain a much better mask (see “ReCLIP++” in the first column). In Fig. 4 (b), we aim to show that the space-preference bias is effectively modeled and rectified. By comparing the results of the original CLIP (see the dashed boxes of “Query” mask) and ReCLIP++ (see the dashed boxes of the last two columns), we find that the segmentation performance in the boundary areas is largely improved. Moreover, the segmentation masks after the rectification stage (ReCLIP++ (rec.) in the penultimate column) appears similar to those after the distillation stage (ReCLIP++ in the last column), which illustrates that the space-preference bias is effectively modeled and rectified in the rectification stage.
@@
(2) Quantitative Results of Bias Rectification. We further design two metrics to quantitatively evaluate the rectification effect of the rectification stage with respect to class-preference and space-preference bias respectively. For class-preference bias, we first compute the confusion matrix for all classes in the dataset. Given a class $i$, we take the $i$-th row of the confusion matrix. We perform a subtraction, between the $i$-th element and the maximum value among other positions of this row, as the score for class $i$. Then we average the scores for all classes. Higher averaged score reflects lower bias and better rectification effect. For space-preference bias, we compute the slopes of the curve between the distance and mIoU at different distances (see Fig. 1 and Appendix for how to draw the curve) which roughly denote $\triangle\text{mIoU}/\triangle\text{Distance}\in(-\infty,0]$. We average the slopes at different distances. Obviously, a lower space-preference bias or a better rectification of space-preference bias should bring a higher score.
@@
We show results on PASCAL Context for class-preference bias, and results on PASCAL VOC for space-preference bias, as there exhibits more obvious class-preference bias on PASCAL Context while more obvious space-preference bias on PASCAL VOC. According to results shown in VII, both ReCLIP (rec.) and ReCLIP++ (rec.) can obviously rectify the class-preference and space-preference bias, compared to MaskCLIP which represents the original segmentation ability of CLIP. Compared to ReCLIP (rec.), ReCLIP++ (rec.) exhibits stronger rectification effect on both biases. Moreover, the quantitative scores with respect to both biases are similar between MaskCLIP+ and MaskCLIP, showing that the bias cannot be effectively rectified by a simple distillation.
@@
TABLE VII: Quantitative Results of Bias Rectification. We design two metrics to evaluate the rectification effect on class-preference and space-preference bias respectively. Higher scores indicate lower bias or stronger rectification effect.
Method Class-preference Rec. Space-preference Rec.
@@
ReCLIP++ (rec.) 0.37 -0.29
Effect of different distillation loss terms. We conduct experiments on PASCAL VOC to validate the effect of each loss term of our distillation framework, including the mask-guided, feature-guided and text-guided loss terms. From Table VIII, all the loss terms contribute to a better performance.
TABLE VIII: Ablation results of distillation loss on PASCAL VOC. Results verify that each loss term contributes to a better performance.
Mask-Guided Feature-Guided Text-Guided mIoU
[…]
TABLE IX: Effect of components for distillation. Results show that it is better to inherit and freeze the components, including query segmentation head, bias logits and the Rectified Mask Generation Module, from the rectification stage than learning the original type of classification head.
Ori. Head Query Bias Logits RMG mIoU
[…]
Effect of components for distillation. As illustrated in Fig. 2 (b), instead of learning a classification head at the distillation stage, the student network (i.e., DeepLab V2 [67]) inherits and freezes the Query segmentation head, bias logits and the Rectified Mask Generation Module from our rectification stage. In Table IX, we conduct experiments to compare learning original type of head (“Ori. Head”) of DeepLab V2 with directly inheriting components including Query segmentation head (“Query”), bias logits (“Bias Logits”) and the Rectified Mask Generation Module (“RMG”) from our rectification stage. Relevant experiments are conducted on PASCAL VOC. From the results, we observe that it is better to inherit the components from our rectification stage than learning a classification head.
Qualitative Results. We visualize our segmentation results in Fig. 5. It can be observed that there exist apparent space-preference bias and class-preference bias in the segmentation results of original CLIP (MaskCLIP). With ReCLIP, both biases can be partially removed. Moreover, ReCLIP++ exhibits stronger bias rectification capability compared to ReCLIP, yielding the best segmentation results.
V Conclusion
In this paper, we propose a new framework for language-guided unsupervised semantic segmentation. We observe the bias, including space-preference bias and class-preference bias, exists in CLIP when directly applying CLIP to segmentation task. We propose using Reference text feature to encode class-preference bias and projecting positional embedding to encode space-preference bias independently, and then manage to combine them into a bias logit map by matrix multiplication. By a simple element-wise logit subtraction mechanism, we rectify the bias of CLIP. Then a contrastive loss is imposed to make the bias rectification meaningful and effective. To further improve the segmentation performance, we distill the knowledge from rectified CLIP to advanced segmentation backbone with specifically designed losses. Extensive experiments demonstrate that our method achieve superior segmentation performance compared to previous state-of-the-arts. We hope our work may inspire future research to investigate how to better adapt CLIP to complex visual understanding tasks.
Acknowledgments
This research project is supported by National Natural Science Foundation (NNSF) of China under Grant 92370114 and National Key R&D Program of China under Grant 2022YFB3104900.

### v2

@@ 4.2 Comparison with previous SOTA
Method Pub. Setting VOC Context ADE City Stuff
GroupViT [57] CVPR’22 TL-OVSS $79.7$ $23.4$ $9.2$ $11.1$ $15.3$
TCL [39] CVPR’23 $77.5$ $30.3$ $14.9$ $23.1$ $19.6$
CoCu [62] NeurIPS’24 - - $11.1$ $15.0$ $13.6$
CLIP [19] ICML’21 TF-OVSS $41.8$ $9.2$ $2.1$ $5.5$ $4.4$
MaskCLIP [20] ECCV’22 $70.0^{\dagger}$ $21.7$ $12.2^{\dagger}$ $19.8^{\dagger}$ $13.6$
SCLIP [64] ECCV’24 $76.9^{\dagger}$ $32.0$ $12.0^{\dagger}$ $15.3^{\dagger}$ $20.5$
ClearCLIP [63] ECCV’24 $81.9$ $34.9$ $16.2$ $18.3$ $22.9$
MaskCLIP+ ECCV’22 C-USS $70.0^{\dagger}$ $31.1$ $12.2^{\dagger}$ $25.2^{\dagger}$ $19.5^{\dagger}$
CLIPpy [22] ICCV’23 $54.6$ $13.5$ $22.3$
ReCo [21] NeurIPS’22 $55.2^{\dagger}$ $26.2$ $19.3$ $26.3$
ReCo [21](w/o. CE) NeurIPS’22 $54.8$ $23.1$ $18.2$ $20.8$
CLIP-S4 [23] CVPR’23 $72.0$ $33.6$
ReCLIP (rec.) [24] CVPR’24 $58.5$ $25.8$ $11.1$ $16.2$ $14.6$
ReCLIP [24] CVPR’24 $75.8$ $33.8$ $14.3$ $19.9$ $20.3$
ReCLIP++ Ours $\bf{85.4}$ $\bf{36.1}$ $\bf{16.4}$ $\bf{26.5}$ $\bf{23.8}$
Comparison. The comparisons with previous state-of-the-art methods on five benchmarks are demonstrated in Table 1. From Table 1, we have the following observations: (1) C-USS methods consider the same setting as our method. We observe that ReCLIP++ outperforms previous state-of-the-art C-USS methods obviously, achieving new state-of-the-arts on all five benchmarks. Since ReCo [21] employs a “context elimination” (CE) trick which introduces prior knowledge to assist the training, we also report the results of ReCo by removing this trick (ReCo w/o. CE in Table 1) for a fair comparison. Notably, ReCLIP++ outperforms MaskCLIP+ by $15.4$%, $5.0$%, $4.2$%, $1.3$% and $4.3$% respectively on the five datasets and outperforms CLIP-S4 by $13.4$% and $2.5$% on PASCAL VOC and PASCAL Context respectively. All these results verify the effectiveness of our ReCLIP++ in rectifying the bias of CLIP to assist unsupervised semantic segmentation (2) ReCLIP++ outperforms previous state-of-the-art TF-OVSS methods, implying that the bias of CLIP in complex visual understanding tasks cannot be fully rectified by simply modifying its architecture without any training. (3) TL-OVSS methods consider a different setting and are trained with large-scale image-caption pairs. Strictly speaking, our method cannot be directly compared with those works. However, the superior performance of ReCLIP++ compared to those typical TL-OVSS methods still demonstrates the effectiveness of ReCLIP++.
4.3 Comparison with Conference Version
@@
In Table 2, we verify the effectiveness of each technical improvement of ReCLIP++ beyond the rectification stage of ReCLIP (rec.) [24] (our conference version): 1) we additionally introduce a decoder that takes the rectified logit map and visual feature of CLIP as input and outputs smoother and more contextual masks (denoted as “Decoder”); 2) we design a new strategy to generate a more accurate label hypothesis for each image (denoted as “Label Hypothesis”); 3) we optimize the design of the bias encoding scheme to independently encode the class-preference and space-preference bias (denoted as “Independent”). In Table 2, we observe that when sequentially adding each technical component, the segmentation performance is consistently improved, which verifies the effectiveness of each technical contribution.
@@
Finally, as shown in Table 1, with all these technical improvements, ReCLIP++ remarkably outperforms ReCLIP. Notably, we observe that even without a distillation, ReCLIP++ still outperforms ReCLIP on all benchmarks, which exhibits a stronger bias rectification capability of ReCLIP++ compared with ReCLIP. For example, on PASCAL VOC and Cityscapes, ReCLIP++ outperforms ReCLIP by $9.6\%$ and $6.6\%$ respectively.
Table 2: Technical improvements of ReCLIP++ compared with ReCLIP. The first line is the results of ReCLIP after the rectification stage and the following lines are ablations experiments on the technical improvements of ReCLIP++ beyond ReCLIP. Results show that each newly added technical improvement of ReCLIP++ contributes to better results than ReCLIP.
ReCLIP (rec.) +Decoder +Label Hypothesis +Independent PASCAL VOC PASCAL Context
$\checkmark$ $58.5$ $25.8$
$\checkmark$ $\checkmark$ $75.6$ $28.3$
$\checkmark$ $\checkmark$ $\checkmark$ $77.3$ $31.4$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bf{85.4}$ $\bf{36.1}$
4.4 Ablation study
Table 3: Ablation on whether Query should be learnable. The $\times$ means fixed while $\checkmark$ means learnable.
Query Reference VOC Context Cityscapes
$\checkmark$ $\checkmark$ 11.6 3.6 2.3
$\times$ $\checkmark$ 85.4 36.1 26.5
@@ 4.4 Ablation study
Should Query be learnable? For text inputs $Q$ and $R$, we only make Reference $R$ learnable while keeping Query $Q$ fixed. In this ablation, we study whether the prompt of Query should also be learnable. As shown in Table 3, we conduct experiments with learnable and fixed Query respectively. The numbers show that fixing Query obviously outperforms making Query learnable. The reason is when we make Query learnable, it may also implicitly capture bias, confusing the following bias rectification operations (e.g., bias logit subtraction). Therefore, in our framework, we choose to fix $Q$ to make the bias rectification more effective.
@@
Effect of element-wise subtraction on bias rectification. In order to validate the effect of our element-wise subtraction, we conduct experiments in Table 4. We compare our subtraction mechanism with an alternative solution: instead of subtracting the bias logit map $M_{b}$ from the Query logit map $M_{q}$, we add $M_{b}$ to $M_{q}$. Comparisons shown in the table verify the effectiveness of our subtraction way. As we aim to remove the bias from the original CLIP, we speculate that the subtraction operation may work as a strong prior which regularizes the training to facilitate meaningful and effective bias modeling.
Table 4: Effectiveness of element-wise subtraction. Results verify the effectiveness of our subtraction way.
Subtraction Addition VOC Context Cityscapes
$\times$ $\checkmark$ 77.7 33.3 21.5
$\checkmark$ $\times$ 85.4 36.1 26.5
@@
Sensitivity to $t$ and $r$ in image-level multi-label hypothesis generation. When generating an image-level multi-label hypothesis as discussed in Sec. 3.4, $t$ and $r$ need to be determined. We study the sensitivity of our method to these hyper-parameters. As shown in Fig. 3, we observe that the mIoU of our method firstly increases and then decreases as $t$ or $r$ increases, exhibiting a typical bell curve. Such a curve shows the regularization effect of $t$ and $r$ on the training. By replacing the multi-label hypothesis generation solution in ReCLIP++ with that used in ReCLIP (i.e., “w/o. crop” in the figure), we obtain results that are inferior to ReCLIP++ within a vast range of $t$ or $r$. All the results show that the solution adopted in ReCLIP++ is insensitive to the choices of $t$ and $r$ and is superior to that used in ReCLIP.
Figure 3: Sensitivity to $t$ and $r$ in the image-level multi-label hypothesis generation. (a) Sensitivity to threshold $t$. (b) Sensitivity to $r$.
@@
Effect of Fine-tuning CLIP. We conduct experiments with fine-tuning the image encoder of CLIP in our framework. The results on PASCAL VOC, PASCAL Context, and Cityscapes are shown in Table 5. We find a consistent performance decrease compared with freezing CLIP. This is because fine-tuning CLIP on small-scale downstream datasets may harm the image-text alignment in pre-trained CLIP.
@@
Table 5: Effect of Fine-tuning CLIP. The “Fine-tune CLIP” means we fine-tune the image encoder of CLIP.
Setting VOC Context Cityscapes
Fine-tune CLIP 76.5 33.7 20.4
Frozen CLIP (Ours) 85.4 36.1 26.5
Table 6: Effectiveness of bias rectification. Results further verify that the rectification of both biases contributes to a better segmentation performance.
Class-preference Space-preference VOC Context
$\checkmark$ $\times$ 76.4 34.1
$\times$ $\checkmark$ 76.6 32.4
$\checkmark$ $\checkmark$ 85.4 36.1
@@
Effect of different bias rectification. We conduct experiments to evaluate how the rectification of each bias affects the segmentation performance in Table 6. Technically, when testing the rectification effect of the class-preference bias, we perform average pooling on the positional feature $W_{p}$ along the spatial dimension and expand the pooling result to the original shape of $W_{p}$. Similarly, when testing the rectification effect of the space-preference bias, we perform average pooling on the reference text feature $W_{r}$ along the class dimension and expand the pooling result to the original shape of $W_{r}$. Results on PASCAL VOC and PASCAL Context shown in Table 6 further verify that the rectification of each bias contributes to a better segmentation performance.
@@
Bias Rectification Evaluation. In this ablation, we further evaluate the rectification effect of two biases both qualitatively and quantitatively.
@@
(1) Qualitative evaluation of bias rectification. In Fig. 4 (a), we illustrate how we explicitly model and rectify the class-preference bias. As shown in the first column of Fig. 4 (a), the original CLIP (see “Query” mask) tends to misclassify part of a “person” (see the ground-truth mask denoted as “GT”) into a “boat”. Such a mistake is reflected in the comparison between the logit heatmaps for the “person” channel and the “boat” channel: in the area misclassified as “boat”, the logits for “boat” are relatively higher than those for “person”. In contrast, for the bias logit map, the person-channel logits are quite low, while the boat-channel logits are generally very high, especially for the misclassified area. Consequently, via the proposed element-wise logit subtraction operation, we obtain the rectified logit map (see the last column), where the boat channel is largely suppressed. Finally, we obtain a much better mask (see “ReCLIP++” in the first column). In Fig. 4 (b), we aim to show that the space-preference bias is effectively modeled and rectified. By comparing the results of the original CLIP (see the dashed boxes of “Query” mask) and ReCLIP++ (see the dashed boxes of the last two columns), we find that the segmentation performance in the boundary areas is largely improved.
@@
(2) Quantitative evaluation of bias rectification. We further design two metrics to quantitatively evaluate the rectification results with respect to class-preference and space-preference bias respectively. For class-preference bias, we first compute the confusion matrix for all classes in the dataset. Given a class $i$, we take the $i$-th row of the confusion matrix. We perform a subtraction, between the $i$-th element and the maximum value among other positions of this row, as the score for class $i$. Then we average the scores for all classes. A higher average score reflects lower bias and better rectification results. For space-preference bias, we compute the slopes of the curve between the distance and mIoU at different distances (see Fig. 1 and Appendix for how to draw the curve) which roughly denote $\triangle\text{mIoU}/\triangle\text{Distance}$. We average the slopes at different distances. Obviously, a lower space-preference bias or a better rectification of space-preference bias should bring a higher score.
@@
We show results on PASCAL Context for class-preference bias, and results on PASCAL VOC for space-preference bias, as there exhibits more obvious class-preference bias on PASCAL Context while more obvious space-preference bias on PASCAL VOC. According to results shown in Table 7, both ReCLIP (rec.) and ReCLIP++ can obviously rectify the class-preference and space-preference bias, compared with MaskCLIP which represents the original segmentation ability of CLIP. Compared with ReCLIP (rec.), ReCLIP++ exhibits stronger rectification results on both biases.
@@
Table 7: Quantitative Evaluation of Bias Rectification. We design two metrics to evaluate the rectification effect on class-preference and space-preference bias respectively. Higher scores indicate lower bias or stronger rectification effect.
Method Class-preference Space-preference
@@
ReCLIP++ 0.40 0.05
ReCLIP++ & distill 0.37 -0.17
Why removing the distillation stage? As discussed in Sec. 3.4, in ReCLIP++, we don’t employ the distillation stage as adopted in ReCLIP. We observe that employing an additional distillation stage after ReCLIP++ leads to inferior results, as shown in Table 8. Further, we empirically study the effect of distillation on bias rectification. As shown in Table 7, we observe that instead of mitigating the bias, further distillation after ReCLIP++ may enlarge the bias, which has a negative effect on the final segmentation performance. This phenomenon explains why the distillation stage is no longer necessary in ReCLIP++. In ReCLIP++, thanks to $\mathcal{F}_{dec}$, the rectified mask of ReCLIP++ is much smoother and contextual than that of ReCLIP, rendering the negative effect of distillation outweighs the positive effect of distillation and leading to a decrease of segmentation performance.
Table 8: Distillation results of ReCLIP++. Results show that distillation harms the segmentation performance.
Method VOC Context ADE20K
ReCLIP++ $\bf{85.4}$ $\bf{36.1}$ $\bf{16.4}$
Distillation $80.0$ $34.5$ $15.8$
Qualitative Results. We visualize our segmentation results in Fig. 5. It can be observed that there exist apparent space-preference bias and class-preference bias in the segmentation results of the original CLIP (MaskCLIP). With ReCLIP, both biases can be partially removed. Moreover, ReCLIP++ exhibits stronger bias rectification capability compared with ReCLIP, yielding the best segmentation results.
5 Conclusion
In this paper, we propose a new framework for language-guided unsupervised semantic segmentation. We observe the bias, including space-preference bias and class-preference bias, exists in CLIP when directly applying CLIP to segmentation task. We propose using Reference text feature to encode class-preference bias and projecting positional embedding to encode space-preference bias independently, and then manage to combine them into a bias logit map by matrix multiplication. By a simple element-wise logit subtraction mechanism, we rectify the bias of CLIP. Then a contrastive loss is imposed to make the bias rectification meaningful and effective. Extensive experiments demonstrate that our method achieves superior segmentation performance compared with previous state-of-the-arts. We hope our work may inspire future research to investigate how to better adapt CLIP to complex visual understanding tasks.